A Car is travelling on a straight road.  The maximum velocity the car can attain is 24 m/s . The maximum acceleration and decleration it csn attain are 1 & 4  m/s2 respectively . What will be the shortest time the car takes from rest to rest in a distance of 200 m.

Grade:

3 Answers

Khimraj
5 years ago
please check acceleration, decleration, max velocity and distance value and confirm it.
…................................................................................................
Rajdeep
5 years ago
HELLO!
 
Maximum velocity that the car can attain is 24 m/s, so v = 24 m/s.
Maximum acceleration a = 1 m/s2
Maximum retardation = 4 m/s2
Distance given = 200m.
 
Car starts from rest and ends by going to rest.
 
Now, time required by the car to start from rest and accelerate to 24 m/s at a rate of 1m/s2  is:
v = u + at
=> 24 = 0 + 1xt
=> t = 24 seconds.
 
Distance covered in this 24 seconds:
S = ut + ½ at2
=> S = 0 + ½ x 1 x 24 x 24
=> S = 288m, which is coming greater than the given value.
 
So, it is proved that there’s something incorrect in the question.
The question, hence, is invalid.
Advaith
5 years ago
Is it 22.36 (10/5)
.. .. . .. .. ..   .. .. .. ... ... .. .. .. .. .... ... . .. . .. . .. ..  ... .. .. ... .. ... ..  .. ..  ........ ..  ... ... .... ...... ... ....  ..  .. .. .. .. ..  .. .